OpenStudy (anonymous):

2. Complete the following ordered pairs for the equation. 6x + -5y = 20 (-5, _) (_, -4) (_, 2)

OpenStudy (imstuck):

Plug in each of those x values and in the blank in each set of parenthesis goes the y value that results.

OpenStudy (anonymous):

plug in the ordered pairs for x,y 6(-5)+(-5y)=20 6x+(-5(-4))=20 6x+(-5(2))=20